OpenStudy (anonymous):

find arc length: y= x^4/8 +1/4x^2 . 12

OpenStudy (blockcolder):

The integral that needs to be found is \[\int\limits_{1}^{2}\sqrt{1+(x^3/2-1/2x^3)}dx\]. Simplify the integrand first before evaluating it.

OpenStudy (anonymous):

\[ 1 + y'^2 =\frac{\left(x^6+1\right)^2}{4 x^6}\\ \sqrt{1 + y'^2} =\frac{\left(x^6+1\right)}{2 x^3}= \\ \frac 1 2x^2 + \frac 1 {2 x^3} \] Integrate from 1 to 2 now

OpenStudy (anonymous):

i dont get it... im getting different stuff:

OpenStudy (anonymous):